# Origen: Zero-Shot 3D Orientation Grounding in Text-to-Image Generation

### v1

@@ C.1 Non-Convergence of Direct Time-Rescaled SDE
Now, consider the reward-guided Langevin SDE in Eq. 4 :
@@
which has the stationary distribution $q^{*}(\mathbf{x})$ given by Eq. 2 . By comparing with the general form, we identify:
@@
$\mathcal{L}^{*}q^{*}(\mathbf{x})$ $=-\nabla\cdot\Big{[}\Big{(}-\frac{1}{2}\mathbf{x}+\eta\,\nabla\hat{\mathcal{R}}(\mathbf{x})\Big{)}\,q^{*}(\mathbf{x})\Big{]}$ (22)
$\quad+\frac{1}{2}\nabla^{2}\Big{[}q^{*}(\mathbf{x})\Big{]}$ (23)
@@
~$\tilde{\mathcal{L}}^{*}q^{*}(\tilde{\mathbf{x}})$ $=-\nabla\cdot\Big{[}\mathcal{G}(\tilde{\mathbf{x}})\Big{(}-\frac{1}{2}\tilde{\mathbf{x}}+\eta\,\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}})\Big{)}q^{*}(\tilde{\mathbf{x}})\Big{]}$
$\quad\quad+\frac{1}{2}\nabla^{2}\Big{[}\mathcal{G}(\tilde{\mathbf{x}})q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=-\nabla\cdot\Big{[}a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})$
$\quad\quad+\mathcal{G}(\tilde{\mathbf{x}})\,\nabla q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=-\nabla\cdot\Big{[}a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})$
$\quad\quad+2a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}\neq 0.$ (28)
which is nonzero in general. Therefore, $q^{*}(\mathbf{x})$ is not annihilated by $\tilde{\mathcal{L}}^{*}$, implying that the time-rescaled SDE does not converge to the desired target distribution in Eq. 2 .
@@ C.2 Time-Rescaled SDE with Correction Drift Term
$d\tilde{\mathbf{x}}_{\tau}=\,$ $\mathcal{G}(\tilde{\mathbf{x}}_{\tau})\left(-\frac{1}{2}\tilde{\mathbf{x}}_{\tau}+\eta\,\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau})\right)d\tau$
$\quad+\frac{1}{2}\nabla\mathcal{G}(\tilde{\mathbf{x}}_{\tau})d\tau$
$\quad+\sqrt{\mathcal{G}(\tilde{\mathbf{x}}_{\tau})}\,d\tilde{W}_{\tau}.$ (29)
@@
~$\tilde{\mathcal{L}}^{*}_{\text{corr}}q^{*}(\tilde{\mathbf{x}})$ $=-\nabla\cdot\Big{[}\mathcal{G}(\tilde{\mathbf{x}})\Big{(}-\frac{1}{2}\tilde{\mathbf{x}}+\eta\,\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}})\Big{)}q^{*}(\tilde{\mathbf{x}})$
$\quad\quad+\frac{1}{2}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla^{2}\Big{[}\mathcal{G}(\tilde{\mathbf{x}})q^{*}(\tilde{\mathbf{x}})\Big{]}$
@@
To verify the effectiveness of our time-rescaling approach, we analyzed the average number of iterations required to reach the desired reward level during reward-guided Langevin dynamics on MS-COCO-Single. Without time rescaling, our method required an average of 14.18 iterations to reach the target reward, whereas applying reward-adaptive time rescaling reduced this to 12.88 iterations, demonstrating improved convergence speed. Notably, the computational overhead introduced by adaptive rescaling is negligible, as the term $\nabla\log\mathcal{G}(\hat{\mathcal{R}}_{i})$ in Alg. 1  can be efficiently computed via the chain rule, using the precomputed reward gradient ^$\nabla\hat{\mathcal{R}}_{i}$.
D Setup for the Toy Experiment
$\mathcal{R}(x,y)$ $=\exp\left(-\frac{(x-4)^{2}+y^{2}}{2}\right)$
$\quad+0.1\cdot\exp\left(-\frac{(x+4)^{2}+y^{2}}{2}\right)-1.$ (31)
In Fig. 2 , all methods use the same total number of sampling steps.
@@ 25 Zebra
Table A: Selected object classes in our dataset.
E General Orientation Controllability
Table B: Quantitative Results on General Orientation Controllability. Origen maintains high accuracy even when evaluated on a more general set of samples.
Model Orientation Alignment Text Alignment
@@ E General Orientation Controllability
Origen (Ours) 0.777 24.96 0.575 12.46 0.969 1.52 0.263 0.710 0.219
We provide our extensive evaluation on general curated MS-COCO-Single dataset. As discussed in Sec. 4.1, we filter out non-clear object classes and image captions, but we do not apply filtering on the front range and do not fix the polar and rotation angles. Upon this, we mix-match object classes and grounding orientations, forming a dataset consisting of 25 object classes, each with 40 samples, totaling 1K samples. We evaluated on same metrics as in Sec. 4.2 , including polar and rotation accuracy, within a tolerance of ±5.0° as well as their absolute errors, following Wang et al. [67]. As shown in Tab. LABEL:supptable:overall_table, Origen generalizes well on diverse orientation conditions. This demonstrates the robustness of our approach, maintaining high orientation alignment performance even when evaluated on a more general set of samples.
F Text-to-Image Alignment Results on MS-COCO-NView
Table C: Quantitative comparisons on 3D orientation grounded image generation on MS-COCO-NView dataset. Origen denotes ours without reward-adaptive time rescaling, where Origen represents our full method. Best and second-best results are highlighted in bold and underlined, respectively.
Id Model 3-view Alignment 4-view Alignment
@@ F Text-to-Image Alignment Results on MS-COCO-NView
4 C3DW [11] 0.504 53.53 0.187 0.334 0.188 - - - - -
5 Zero-1-to-3 [46] 0.366 68.70 0.266 0.646 0.210 0.321 75.10 0.264 0.642 0.209
@@
In Tab. C, we report the additional quantitative results on text-to-image alignment on the MS-COCO-NView benchmark. As noted in Sec. 4.4 , Origen achieves comparable text alignment results compared to the vanilla T2I models, while achieving significantly more accurate orientation grounding.
G User Study Examples
@@
Figure A: Screenshots of our user study.
@@
In this section, we provide details of the user study. To assess user preferences, we conducted an evaluation comparing images generated by Zero-1-to-3 [46], C3DW [11], and Origen using MS-COCO-Single as the benchmark. The study was conducted on Amazon Mechanical Turk (AMT). For each object class, one sample was randomly selected, resulting in a total of 25 questions. The orientations used to generate the images were intuitively visualized and presented alongside their corresponding prompts. As shown in Fig. A, participants were asked to respond to the following question: “Considering the orientation and prompt conditions, which image most closely follows ALL the conditions below?" Each user study session included 25 test samples along with 5 vigilance tests, which were randomly interspersed. The final results were derived from responses of valid participants who correctly answered at least three vigilance tests, leading to a total of 55 valid participants out of 100. No additional eligibility restrictions were imposed.
H Implementation Details
Figure B: Plot of our monitor function $\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}))$.
About Origen. We employed FLUX-Schnell [34] as our one-step T2I generative model and OrientAnything (ViT-L) [67] to measure the Orientation Grounding Reward, as detailed in Sec. 3.2. For all experiments, we set $\gamma=0.3$ and $\eta=1$ in Alg. 1, as this configuration provides a favorable balance between image quality and computational cost when evaluating with 50 NFEs. As described in Sec. 3.4, we additionally visualize our monitor function $\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}))$ from Eq. 9 in Fig. B, which demonstrates how the function adaptively scales the step size–assigning smaller steps in high-reward regions and larger steps in low-reward regions–thereby improving convergence speed.
About Zero-1-to-3 [46]. For the Zero-1-to-3 [46] baseline, we generated the base image using FLUX-Schnell with 4 steps and encouraged the model to generate front-facing objects by adding the “facing front" prompt. The model was then conditioned on the target azimuth while keeping the polar angle fixed at 0° to generate novel foreground views, which were later composited with the background. The foreground was segmented using SAM[29], and missing background pixels were inpainted using LaMa [61] before composition.
About C3DW [11]. For C3DW [11], we utilized the orientation & illumination model checkpoint provided in the official implementation. This model is only capable of controlling azimuth angles for half-front views and was trained with scalar orientation values ranging from 0.0 to 0.5, where 0.0 corresponds to 90°, and 0.5 corresponds to -90°, with intermediate orientations obtained through linear interpolation. Using this mapping, we converted ground truth (GT) orientations into the corresponding input values for the model.
Handling Back-Facing Generation. We empirically observed that our base model struggled to generate back-facing images when the prompt did not explicitly contain view information. We hypothesized that this issue arises because high-reward samples lie within an extremely sparse region of the conditional probability space, making them inherently challenging to sample, even when employing our proposed approach. To address this issue, we explicitly added the phrase “facing back" in facing back cases (i.e., where 90 $<$ $\phi_{i}^{\textit{az}}$ $<$ 270) and optimized the noise accordingly. With this simple adjustment, the model effectively generated both facing front and back images while maintaining high-quality outputs.

### v2

@@ C.1 Non-Convergence of Direct Time-Rescaled SDE
Now, consider the reward-guided Langevin SDE in Eq. 4 in the main paper:
@@
which has the stationary distribution $q^{*}(\mathbf{x})$ given by Eq. 2 in the main paper. By comparing with the general form, we identify:
@@
$\mathcal{L}^{*}q^{*}(\mathbf{x})$ $=-\nabla\cdot\Big{[}\Big{(}-\frac{1}{2}\mathbf{x}+\eta\,\nabla\hat{\mathcal{R}}(\mathbf{x})\Big{)}\,q^{*}(\mathbf{x})\Big{]}+\frac{1}{2}\nabla^{2}\Big{[}q^{*}(\mathbf{x})\Big{]}$
@@
~$\tilde{\mathcal{L}}^{*}q^{*}(\tilde{\mathbf{x}})$ $=-\nabla\cdot\Big{[}\mathcal{G}(\tilde{\mathbf{x}})\Big{(}-\frac{1}{2}\tilde{\mathbf{x}}+\eta\,\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}})\Big{)}q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla^{2}\Big{[}\mathcal{G}(\tilde{\mathbf{x}})q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=-\nabla\cdot\Big{[}a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})+\mathcal{G}(\tilde{\mathbf{x}})\,\nabla q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=-\nabla\cdot\Big{[}a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})+2a(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}$
$=\frac{1}{2}\nabla\cdot\Big{[}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}$
$\neq 0,$ (26)
where we used the fact that $\mathcal{G}(\mathbf{x})\,\nabla q^{*}(\mathbf{x})=2a(\mathbf{x})\,q^{*}(\mathbf{x})$, which follows from the identity $\nabla\log q^{*}(\mathbf{x})=-\mathbf{x}+2\eta\,\nabla\hat{\mathcal{R}}(\mathbf{x})$. Therefore, $q^{*}(\mathbf{x})$ is not annihilated by $\tilde{\mathcal{L}}^{*}$, implying that the time-rescaled SDE does not converge to the desired target distribution in Eq. 2 in the main paper.
@@ C.2 Time-Rescaled SDE with Correction Drift Term
$d\tilde{\mathbf{x}}_{\tau}=\,$ $\mathcal{G}(\tilde{\mathbf{x}}_{\tau})\left(-\frac{1}{2}\tilde{\mathbf{x}}_{\tau}+\eta\,\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau})\right)d\tau+\frac{1}{2}\nabla\mathcal{G}(\tilde{\mathbf{x}}_{\tau})d\tau+\sqrt{\mathcal{G}(\tilde{\mathbf{x}}_{\tau})}\,d\tilde{W}_{\tau}.$ (27)
@@
~$\tilde{\mathcal{L}}^{*}_{\text{corr}}q^{*}(\tilde{\mathbf{x}})$ $=-\nabla\cdot\Big{[}\mathcal{G}(\tilde{\mathbf{x}})\Big{(}-\frac{1}{2}\tilde{\mathbf{x}}+\eta\,\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}})\Big{)}q^{*}(\tilde{\mathbf{x}})+\frac{1}{2}\nabla\mathcal{G}(\tilde{\mathbf{x}})\,q^{*}(\tilde{\mathbf{x}})\Big{]}+\frac{1}{2}\nabla^{2}\Big{[}\mathcal{G}(\tilde{\mathbf{x}})q^{*}(\tilde{\mathbf{x}})\Big{]}$
@@
Table 3: Convergence speed to the desired success criterion. Origen denotes ours without reward-adaptive time rescaling.
Method Origen Origen
Mean NFE $\downarrow$ 14.2 12.8
Inference time (s) $\downarrow$ 14.9 13.5
@@
To verify the effectiveness of our time-rescaling approach, we analyzed the average number of iterations required to reach the desired success criterion during reward-guided Langevin dynamics on ORIBENCH-Single dataset. As shown in Tab. 3, Origen requires an average of 14.2 iterations to reach the desired reward level, whereas Origen reduced this to 12.8 iterations, demonstrating improved convergence speed without image quality degradation. Notably, the computational overhead introduced by adaptive rescaling is negligible, as the term $\nabla\log\mathcal{G}(\hat{\mathcal{R}}_{i})$ in Alg. 1 in the main paper can be efficiently computed via the chain rule, using the precomputed reward gradient ^$\nabla\hat{\mathcal{R}}_{i}$.
@@ 25 Zebra
Table 4: Selected object classes in our dataset.
Appendix D Setup for the Toy Experiment
$\mathcal{R}(x,y)$ $=\exp\left(-\frac{(x-4)^{2}+y^{2}}{2}\right)+0.1\cdot\exp\left(-\frac{(x+4)^{2}+y^{2}}{2}\right)-1.$ (29)
In Fig. 2 in the main paper, all methods use the same total number of sampling steps.
Appendix E Selected Classes
When constructing the ORIBENCH benchmark, we selected a total of 25 classes from the 80 object classes in MS-COCO validation set [23]. We excluded classes for which distinguishing the front and back is difficult or where defining a canonical orientation is ambiguous. The remaining 25 classes were chosen based on their clear orientation cues. The detailed list of selected classes is provided in Tab. 4.
Appendix F Implementation Details
Figure 5: Plot of our monitor function $\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}))$.
About Origen.
We employed FLUX-Schnell [21] as our one-step T2I generative model and OrientAnything (ViT-L) [20] to measure the Orientation Grounding Reward, as detailed in Sec. 3.2. For all experiments, we set $\eta=0.8$ in Alg. 1, and used $\gamma=0.3$ for ORIBENCH-Single and $\gamma=0.2$ for ORIBENCH-Multi, as this configuration provides a favorable balance between image quality and computational cost when evaluating with 50 NFEs. As described in Sec. 3.4, we additionally visualize our monitor function $\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}))$ from Eq. 9 in Fig. 5, which demonstrates how the function adaptively scales the step size–assigning smaller steps in high-reward regions and larger steps in low-reward regions–thereby improving convergence speed.
About Guided Generation Methods.
For ReNO [22], we employed FLUX-Schnell [21] as the base one-step text-to-image generative model and OrientAnything (ViT-L) [20] as in ours. For multi-step guidance methods, we used FLUX-Dev [21] as multi-step text-to-image generative model. We set the gradient weight hyperparameter to $0.3$ for ORIBENCH-Single and 0.2 for ORIBENCH-Multi for all methods. In the case of ReNO [22], we use norm-based regularization weight of $0.01$. To ensure a fair comparison, we matched the number of function evaluations (NFEs) for all training-free one-step, multi-step guidance methods to 50.
About Zero-1-to-3 [17].
For the Zero-1-to-3 [17] baseline, we generated the base image using FLUX-Schnell (512$\times$512) with 4 steps and encouraged the model to generate front-facing objects by adding the “facing front" prompt. The model was then conditioned on the target azimuth while keeping the polar angle fixed at 0° to generate novel foreground views, which were later composited with the background. The foreground was segmented using SAM [76], and missing background pixels were inpainted using LaMa [77] before composition.
About C3DW [18].
For C3DW [18], we utilized the orientation & illumination model checkpoint provided in the official implementation. This model is only capable of controlling azimuth angles for half-front views and was trained with scalar orientation values ranging from 0.0 to 0.5, where 0.0 corresponds to 90°, and 0.5 corresponds to -90°, with intermediate orientations obtained through linear interpolation. Using this mapping, we converted ground truth (GT) orientations into the corresponding input values for the model.
Handling Back-Facing Generation.
We empirically observed that our base model struggled to generate back-facing images when the prompt did not explicitly contain view information. We hypothesized that this issue arises because high-reward samples lie within an extremely sparse region of the conditional probability space, making them inherently challenging to sample, even when employing our proposed approach. To address this issue, we explicitly added the phrase “facing back" in facing back cases (i.e., where 90 $<$ $\phi_{i}^{\textit{az}}$ $<$ 270) and optimized the noise accordingly. With this simple adjustment, the model effectively generated both facing front and back images while maintaining high-quality outputs.
Appendix G Additional Results on the Extended ORIBENCH-Single Benchmark
We further evaluate Origen under two additional scenarios that better reflect real-world use cases, by extending the ORIBENCH-Single Benchmark. The first experiment investigates the question: (1) “Can Origen handle more general and complex orientations beyond simple angles?". The second experiment addresses: (2) “Can orientation grounding be achieved simply by prompting a text-to-image generative model?". For the first scenario, we evaluate Origen and other guided generation methods on an extended benchmark covering all three orientation components—azimuth, polar, and rotation without filtering the front range of azimuths. For the second scenario, we assess the capability of text-to-image models to perform orientation grounding for four primitive directions (front, left, right, back), where the orientation condition is provided via the input text prompt.
Table 5: Quantitative comparisons on General Orientation Controllability. Origen maintains high accuracy even when evaluated on a more general set of samples. Best and second-best results are highlighted in bold and underlined, respectively.
Id Model Orientation Alignment Text-to-Image Alignment
@@ Appendix G Additional Results on the Extended ORIBENCH-Single Benchmark
1 DPS [32] 0.432 43.89 0.511 13.07 0.969 1.66 0.259 0.704 0.234
2 MPGD [33] 0.398 48.09 0.485 12.25 0.967 1.79 0.258 0.707 0.234
3 FreeDoM [35] 0.470 38.09 0.554 12.71 0.971 1.42 0.261 0.709 0.229
4 ReNO [22] 0.586 41.41 0.502 14.37 0.958 2.69 0.253 0.676 0.216
5 Origen (Ours) 0.777 24.96 0.575 12.46 0.969 1.52 0.263 0.710 0.219
Table 6: Quantitative comparisons on 3D orientation grounded image generation for four primitive views. Best and second-best results are highlighted in bold and underlined, respectively.
3-View (front, left, right) 4-View (front, left, right, back)
Id Model Orientation Alignment Text-to-Image Alignment Orientation Alignment Text-to-Image Alignment
@@
(2) Fine-tuned Orientation-to-Image Model
4 Zero-1-to-3 [17] 0.366 59.03 0.266 0.646 0.210 0.321 75.10 0.264 0.642 0.209
5 C3DW [18] 0.504 64.77 0.187 0.334 0.188
(3) Guided-Generation Methods with One-Step Model
@@
Orientation DPS [32] MPGD [33] FreeDoM [35] ReNO [22] Origen (Ours)
Figure 6: Qualitative comparisons on generally extended ORIBENCH-Single benchmark. Compared to other training-free approaches [32, 33, 35, 22], Origen generates the best aligned images with the given orientation grounding conditions.
Figure 7: Qualitative comparisons on extend ORIBENCH-Single benchmark for four primitive orientations.
G.1 Orientation Grounding in more general and complex scenario.
We provide our extensive evaluation on general curated ORIBENCH-Single dataset. As discussed in Sec. 4.1, we filter out non-clear object classes and image captions, but we do not apply filtering on the front range and do not fix the polar and rotation angles. Upon this, we mix-match object classes and grounding orientations, forming a dataset consisting of 25 object classes, each with 40 samples, totaling 1K samples. We evaluated on same metrics as in Sec. 4.2 in the main paper, including polar and rotation accuracy, within a tolerance of ±5.0° as well as their absolute errors, following Wang et al. [20]. As shown in Tab. 5, Origen generalizes well on diverse orientation conditions, outperforming all other training-free guidance methods. We also present qualitative comparisons for other training-free guidance methods in Fig. 6 . Origen shows more consistent and accurate orientation alignment compared to other training-free approaches. This demonstrates the robustness of our approach, maintaining high orientation alignment performance even when evaluated on a more general and complex set of samples.
G.2 Text-to-Image Models.
As baselines, we consider several one-step T2I generative models: SD-Turbo [24], SDXL-Turbo [25], and FLUX-Schnell [21]. In particular, we appended a phrase that specifies the object orientation at the end of each caption in ORIBENCH-Single dataset. For more comprehensive comparisons, we also included the fine-tuned models (C3DW [18] and Zero-1-to-3 [17]) considered in Sec. 4.5 in this experiment as well. As shown in Tab. 6, Origen significantly outperforms all baseline models in orientation alignment. Note that, although FLUX-Schnell [21] achieves the highest alignment among the vanilla T2I models, Origen surpasses it by more than 2.5 times in the 3-view alignment setting (82.4% vs. 31.2%) and more than 2 times in the 4-view alignment setting (86.6% vs. 42.4%). This substantial margin highlights the inherent ambiguity and lack of precise control in vanilla T2I approaches, as orientation information embedded within textual descriptions is less explicit and reliable compared to direct orientation guidance. We further demonstrate the advantage of Origen through qualitative comparisons in Fig. 7. Vanilla T2I models frequently fail to adhere strictly to the desired orientation, even with directional phrases in text prompts. In contrast, Origen consistently generates images accurately aligned with the specified orientations.
@@
Figure 8: Screenshots of our user study.
Appendix H User Study Examples
@@
In this section, we provide details of the user study. To assess user preferences, we conducted an evaluation comparing images generated by Zero-1-to-3 [17], C3DW [18], and Origen using ORIBENCH-Single as the benchmark. The study was conducted on Amazon Mechanical Turk (AMT). For each object class, one sample was randomly selected, resulting in a total of 25 questions. The orientations used to generate the images were intuitively visualized and presented alongside their corresponding prompts. As shown in Fig. 8, participants were asked to respond to the following question: “Considering the orientation and prompt conditions, which image most closely follows ALL the conditions below?”. Each user study session included 25 test samples along with 5 vigilance tests, which were randomly interspersed. The final results were derived from responses of valid participants who correctly answered at least three vigilance tests, leading to a total of 55 valid participants out of 100. No additional eligibility restrictions were imposed.
Table 7: Computational cost comparison of orientation grounding methods on single NVIDIA A100 GPU.
Method NFEs Time per Iter (Total) VRAM Img Size SD-Turbo [24] 1 0.08s (0.08s) 4GB $512\times 512$ SDXL-Turbo [25] 1 0.30s (0.30s) 15GB $1024\times 1024$ FLUX-Schnell [21] 1 1.01s (1.01s) 32GB $512\times 512$ C3DW [18] 20 6.09s (0.30s) 6GB $768\times 768$ Zero-1-to-3 [17] 50 8.19s (0.16s) 46GB $256\times 256$ DPS [32] 50 53.2s (1.06s) 43GB $512\times 512$ MPGD [33] 50 39.3s (0.79s) 36GB $512\times 512$ FreeDoM [35] 50 52.9s (1.06s) 43GB $512\times 512$ ReNO [22] 50 54.5s (1.09s) 43GB $512\times 512$ Origen (Ours) 50 52.7s (1.05s) 43GB $512\times 512$
Appendix I Computational Costs
In Tab. 7, we compare the computational costs of all baselines and Origen in terms of inference time and GPU memory consumption. Our base model (FLUX-Schnell [21]) and all guided generation methods were evaluated for generating 512$\times$512 images, while other methods followed their default settings. Guided generation methods are generally slower due to the need for reward backpropagation. All methods were evaluated under a setup of 50 Number of Function Evaluations (NFEs). Further analysis demonstrating the fast convergence of our reward-guided Langevin Dynamics is provided in Appendix C.2, suggesting that the computational cost can be partially reduced by using fewer NFEs. All measurements were conducted on a single A100 GPU with 80GB of memory.